a motorboat takes 3 hours to travel 180 mi going upstream. the return trip takes 2 hours going downstream. what is the rate of the boat in still water and what is the rate of the current?
Answer by lwsshak3(11628)

a motorboat takes 3 hours to travel 180 mi going upstream. the return trip takes 2 hours going downstream. what is the rate of the boat in still water and what is the rate of the current?
***
let x=
x+c=rate of boat downstream
x-c=rate of boat upstream
distance=travel time*rate
..
3(x-c)=180
2(x+c)=180
..
3x-3c=180
2x+2c=180
..
6x-6c=360(mult. by 2)
6x+6c=540(mult. by 3)
add
12x=900
x=75
2c=180-2x=30
c=15
rate of boat in still water=75 mph
rate of current=15 mph